# Structure of activity in multiregion recurrent neural networks

### v1

@@ II Multiregion network model
This connectivity allows selective communication of activity between regions in a manner controlled by the geometry encoded in $T^{\mu\nu\rho}$. We first explain this heuristically. If $\phi^{\nu}_{i}$ is an activity pattern in region $\nu$, the activity communicated to region $\mu$ is proportional to the projection $\sum_{i}{n}^{\mu\nu}_{i}\phi^{\nu}_{i}/N$. For a generic pattern $\phi^{\nu}_{i}$ (e.g., induced by the disordered connectivity $\chi^{\nu}_{ij}$), this projection is of order $1/\sqrt{N}$, and thus vanishes for $N\rightarrow\infty$. However, if $\phi_{i}^{\nu}$ has a component along $n_{i}^{\mu\nu}$, this projection is order-one. For this to be the case, there must be a region $\rho$ such that $m^{\nu\rho}_{i}$, which delivers input to region $\nu$, has a component along $n_{i}^{\mu\nu}$. This component is $T^{\mu\nu\rho}$. It follows from this argument that high-dimensional chaotic activity is unable to propagate between regions as $N\rightarrow\infty$.
III Spectral analysis
Figure 2: Tensor diagrams illustrating the relationships between (a) matrices $\bm{X}\bm{Y}$ and $\bm{Y}\bm{X}$, and (b) the matrices with more complicated indices relevant to Sec. III. In both (a) and (b), the tensors on the left and right have the same eigenvalues, up to zeros, with respect to the left-right bipartition of the indices. Dangling legs are indices of the output tensor, connected legs are summed over, and $\bullet$ is the Kronecker delta. Dashed lines are summed over on the left and split open on the right; solid lines are dangling on the left and joined on the right.
Before proceeding to the mean-field analysis, we describe the spectrum of $J^{\mu\nu}_{ij}$, which controls the linearized dynamics about the trivial fixed point, $x^{\mu}_{i}=0$. Rather than as a fourth-order tensor, $J^{\mu\nu}_{ij}$ can be interpreted as an $RN$-by-$RN$ matrix with respect to the two “superindices” $(\mu,i)$ and $(\nu,j)$. In other words, the multiregion network can be interpreted as a single large network. This $RN$-by-$RN$ matrix has a bulk spectrum resulting from the random matrices $\chi_{ij}^{\mu}$, whose density in the complex plane, for $N\rightarrow\infty$, is a superposition of $R$ uniform disks of radii $g^{\mu}$.
We denote the low-rank part of the connectivity by ${L^{\mu\nu}_{ij}=m^{\mu\nu}_{i}n^{\mu\nu}_{j}/N}$. This term has up to $R^{2}$ nonzero eigenvalues, which do not interact with the bulk as $N\rightarrow\infty$; they either are outliers or are swallowed by the bulk. To determine these eigenvalues, we seek an $R^{2}$-by-$R^{2}$ matrix whose spectrum coincides with that of $L^{\mu\nu}_{ij}$ as $N\rightarrow\infty$. Such a matrix can be found using the property that the matrices $\bm{X}\bm{Y}$ and $\bm{Y}\bm{X}$ have the same spectra up to zeros (Fig. 2(a)). First, we express the low-rank component as a product of matrices,
$L^{\alpha\beta}_{ij}=\sum_{\mu,\nu}\frac{\delta^{\mu\alpha}m_{i}^{\mu\nu}}{\sqrt{N}}\frac{\delta^{\nu\beta}n_{j}^{\mu\nu}}{\sqrt{N}},$ (7)
contracting over the superindex $(\mu,\nu)$ to form a matrix with superindices $(\alpha,i)$ and $(\beta,j)$. The same eigenvalues, up to zeros, are obtained by contracting instead over the superindex $(\alpha,i)=(\beta,j)$, resulting in a matrix with superindices $(\mu,\nu)$ and $(\rho,\sigma)$. This $R^{2}$-by-$R^{2}$ matrix is given by
$\hat{T}^{\mu\nu,\rho\sigma}$ $=\delta^{\nu\rho}\frac{1}{N}\sum_{i}n_{i}^{\mu\nu}m_{i}^{\nu\rho}$
$=\delta^{\nu\rho}T^{\mu\nu\sigma},$ (8)
where the limit $N\rightarrow\infty$ was taken in the second line. This matrix can also be derived in a single step using a tensor diagram (Fig. 2(b)).
When all eigenvalues of this matrix and the bulk have real parts less than unity, the trivial fixed point of the network is stable, leading to quiescent behavior. If any eigenvalue exceeds this threshold, the network exhibits nontrivial activity described by the DMFT.
IV Dynamical mean-field theory (DMFT)
In the mean-field description of the network, the order parameters comprise the $R$ region-specific two-point functions,
$\Delta^{\mu}(t,t^{\prime})$ $=\left\langle x^{\mu}_{i}(t)x^{\mu}_{i}(t^{\prime})\right\rangle_{\bm{J}},$ (9a)
$C^{\mu}(t,t^{\prime})$ $=\left\langle\phi^{\mu}_{i}(t)\phi^{\mu}_{i}(t^{\prime})\right\rangle_{\bm{J}},$ (9b)
together with $R^{2}$ additional parameters describing pairwise communication between regions,
[…]
which is the activity in region $\nu$ that gets communicated to region $\mu$ (after being low-pass filtered). We refer to these variables as currents, borrowing the terminology of Perich et al. [9]. In the mean-field picture, these currents interact according to
$(1+\partial_{t})S^{\mu\nu}(t)$ $=H^{\mu\nu}(t),\text{ where}$ (11a)
$H^{\mu\nu}(t)$ $=\psi^{\nu}(t)\sum_{\rho}{T}^{\mu\nu\rho}S^{\nu\rho}(t),$ (11b)
which is a matrix dynamics shaped by a third-order tensor. Here, $\psi^{\nu}(t)=\left\langle\phi^{\prime}(x^{\nu}_{i}(t))\right\rangle_{\bm{J}}$ is the average gain of neurons in region $\nu$ at time $t$. Due to the Gaussianity of the preactivations, this can be expressed as $\psi^{\nu}(t)=\psi(\Delta^{\nu}(t,t))$ where
$\psi(\Delta)=\left\langle\phi^{\prime}(x)\right\rangle_{x},\>\>\>x\sim\mathcal{N}(0,\Delta).$ (12)
Once $\Delta^{\mu}(t,t)$ has been determined, the dynamics of the currents are given by Eq. IV. Squaring both sides of the network dynamics, Eq. 1, and doing a disorder average gives a dynamic equation for $\Delta^{\mu}(t,t^{\prime})$,
$(1+\partial_{t})(1+\partial_{t^{\prime}})\Delta^{\mu}(t,t^{\prime})=(g^{\mu})^{2}C^{\mu}(t,t^{\prime})\\ +\sum_{\nu,\rho}U^{\mu\nu\rho}H^{\mu\nu}(t)H^{\mu\rho}(t^{\prime}).$ (13)
@@ IV Dynamical mean-field theory (DMFT)
These DMFT equations are closed by expressing $C^{\mu}(t,t^{\prime})$ in terms of $\Delta^{\mu}(t,t^{\prime})$. Again noting the Gaussianity of the preactivations, we have ${C^{\mu}(t,t^{\prime})=C(\Delta(t,t^{\prime}),\Delta(t,t),\Delta(t^{\prime},t^{\prime}))}$, where
$C(\Delta_{12},\Delta_{11},\Delta_{22})=\left\langle\phi(x_{1})\phi(x_{2})\right\rangle_{x_{1},x_{2}},\\ \left(\begin{array}[]{c}x_{1}\\ x_{2}\end{array}\right)\sim\mathcal{N}\left(\left(\begin{array}[]{c}0\\ 0\end{array}\right),\left(\begin{array}[]{cc}\Delta_{11}&\Delta_{12}\\ \Delta_{12}&\Delta_{22}\end{array}\right)\right).$ (14)
In the case of the error-function nonlinearity, the Gaussian integrals in Eqs. 12 and 14 can be evaluated analytically to give
$\psi(\Delta)=\frac{1}{\sqrt{1+{\pi\Delta}/{2}}},$ (15a)
$C(\Delta_{12},\Delta_{11},\Delta_{22})$ (15b)
$\hskip 22.0pt=\frac{2}{\pi}\text{arctan}\frac{\Delta_{12}}{\sqrt{\left(\Delta_{11}+{2}/{\pi}\right)\left(\Delta_{22}+{2}/{\pi}\right)}}.$
Eqs. IV14 comprise a set of deterministic, causal dynamic equations for the $R$ region-specific two-point functions and $R^{2}$ currents. In general, these equations must be solved via numerical integration on a two-dimensional grid. This can be avoided in situations without disorder or with time-translation invariance, both of which we will consider.
For time-independent gains ${\psi^{\nu}(t)=\text{const.}}$. (e.g., at the trivial fixed point), the currents interact via linear dynamics. Note that $\hat{T}^{\mu\nu,\rho\sigma}$, defined in Eq. 8, is the $R^{2}$-by-$R^{2}$ matrix that governs these linear dynamics, i.e.,
$(1+\partial_{t})S^{\mu\nu}(t)\propto\sum_{\rho,\sigma}\hat{T}^{\mu\nu,\rho\sigma}S^{\rho\sigma}(t).$ (16)
For this reason, we expect the spectrum of $\hat{T}^{\mu\nu,\rho\sigma}$ to shape the full dynamics of the currents even in the nonlinear case, in analogy with other scenarios where a local linear instability is informative of global nonlinear behavior [42, 43].
@@
Finally, throughout the rest of the paper, we assume $U^{\mu\nu\rho}=\delta^{\nu\rho}$ for all $\mu$, and focus on the role of $T^{\mu\nu\rho}$. Geometrically, this means that inputs from other regions into region $\mu$ are organized in orthogonal subspaces in region $\mu$.
V Symmetric effective interactions
@@
We first make a common assumption in the study of many-body dynamics, namely, that the interactions are symmetric. Enforcing symmetry in this system is nontrivial, however, because the effective interactions are a third-order tensor $T^{\mu\nu\rho}$. As per the above discussion, a natural choice is to impose symmetry on the matrix $\hat{T}^{\mu\nu,\rho\sigma}$, i.e., $\hat{T}^{\mu\nu,\rho\sigma}=\hat{T}^{\rho\sigma,\mu\nu}$. This constraint reduces the number of free parameters from $\mathcal{O}(R^{3})$ to $\mathcal{O}(R^{2})$ by requiring
$T^{\mu\nu\rho}$ $=\delta^{\mu\rho}c^{\mu\nu},\text{ where }$ (17a)
$c^{\mu\nu}$ $=c^{\nu\mu}.$ (17b)
In contrast, for standard coupling matrices, imposing symmetry does not reduce the scaling of the number of parameters. Note that, even with this constraint, the high-dimensional couplings $J^{\mu\nu}_{ij}$ are not symmetric, both because the matrices $\chi^{\mu}_{ij}$ are asymmetric, and because sampling the vectors $m^{\mu\nu}_{i}$ and $n^{\mu\nu}_{i}$ results in ${m^{\mu\nu}_{i}n^{\mu\nu}_{j}\neq m^{\nu\mu}_{j}n^{\nu\mu}_{i}}$.
The presence of $\delta^{\mu\rho}$ in Eq. 17a implies that each region $\mu$ either interacts directly with itself ($\mu=\nu$) or does so indirectly through an intermediate region, $\nu$ ($\mu\neq\nu$). Moreover, the symmetry of $c^{\mu\nu}$, Eq. 17b, implies that the coupling through which region $\mu$ interacts with itself via region $\nu$ is equivalent to that for region $\nu$ interacting with itself via region $\mu$. In terms of the geometric arrangement of connectivity patterns, these constraints imply that within region $\mu$, the vectors $n^{\mu\mu}_{i}$ and $m^{\mu\mu}_{i}$ have nontrivial overlap; similarly, the vectors $n^{\mu\nu}_{i}$ and $m^{\nu\mu}_{i}$, which define how activity is extracted from region $\nu$ and provided to region $\mu$, also have nontrivial overlap. This latter overlap is the same when $\mu$ and $\nu$ are switched in accordance with the symmetry of $c^{\mu\nu}$. All other overlaps are zero. This geometric arrangement of connectivity patterns underpins the dynamics of subspace-based signal routing, with orthogonality between pairs of vectors replacing the traditional notion of connectivity sparsity. This is illustrated in Fig. 3(a).
To make analytical progress, we further constrain $c^{\mu\nu}$ to have a “rank-one plus diagonal” form with only $\mathcal{O}(R)$ parameters,
@@
where $u^{\mu}$ and $h^{\mu}$ are arbitrary vectors. The expressivity of this parameterization of a symmetric matrix is well known from the factor-analysis model in statistics [44]. This parameterization provides a minimal example in which one has independent control over the strength of direct versus indirect self-connectivity, which are captured by the quantities
$a^{\mu}$ $=(u^{\mu})^{2}+h^{\mu},\text{ and}$ (19a)
$b^{\mu}$ $=(u^{\mu})^{2},$ (19b)
respectively. Finally, the $R^{2}$ eigenvalues of $\hat{T}^{\mu\nu,\rho\sigma}$ are $a^{\mu}$ for all $\mu$ and ±$\pm\sqrt{|b^{\mu}b^{\nu}|}$ for all pairs $\mu\neq\nu$. We will assume that, for each $\mu$, ${\text{max}(a^{\mu},b^{\mu})>1}$.
V.1 Disorder-free case
The lack of disorder simplifies the DMFT because the variance of neurons is due solely to the currents. Mathematically, setting $g^{\mu}=0$ implies that the two-point function dynamics, Eq. 13, can be integrated to give $\Delta^{\mu}(t,t)=\sum_{\nu}(S^{\mu\nu}(t))^{2}$. Generically, symmetric interactions lead to fixed points, and this is the case here. Upon setting $g^{\mu}$ and the time-derivative to zero in the DMFT equations (Eqs. IV14) and further specializing to the parameterization of $T^{\mu\nu\rho}$ given above, fixed points $S^{\mu\nu}_{0}$ of the disorder-free current dynamics satisfy

### v3

@@ 3.3 Functional Significance of Low-Rank Cross-Region Connectivity
This connectivity structure allows selective communication between regions, controlled by the geometry encoded in $T^{\mu\nu\rho}$. To illustrate this mechanism, consider an activity pattern $\phi^{\nu}_{i}$ in region $\nu$. The activity communicated to region $\mu$ is proportional to the projection $N^{-1}\sum_{i=1}^{N}n^{\mu\nu}_{i}\phi^{\nu}_{i}$. For a generic pattern $\phi^{\nu}_{i}$ (e.g., induced by the disordered connectivity $\chi^{\nu}_{ij}$), this projection is of order $1/\sqrt{N}$, vanishing as $N\rightarrow\infty$. However, if $\phi^{\nu}_{i}$ has a component aligned with $\bm{n}^{\mu\nu}$, this projection remains of order unity.
For such alignment to occur, there must exist a region $\rho$ such that $\bm{m}^{\nu\rho}$, which delivers input to region $\nu$, has a component along $\bm{n}^{\mu\nu}$. This component is precisely $T^{\mu\nu\rho}$. Consequently, high-dimensional chaotic activity cannot propagate between regions as $N\rightarrow\infty$, ensuring that only structured, low-dimensional signals are transmitted.
Network variables
$x_{i}^{\mu}(t)$ Preactivation (“membrane potential”) of neuron $i$ in region $\mu$ at time $t$ (Eq. 1)
$\phi_{i}^{\mu}(t)$ Activation (“firing rate”) of neuron $i$ in region $\mu$ at time $t$ (Eq. 1)
Network parameters
$N$ Number of neurons in each region
$R$ Number of regions
$J_{ij}^{\mu\nu}$ Synaptic coupling from neuron $j$ in region $\nu$ to neuron $i$ in region $\mu$ (Eq. 2)
$\chi_{ij}^{\mu}$ Random component of within-region synaptic couplings in region $\mu$ (Eq. 2)
$g^{\mu}$ Standard deviation (times $\sqrt{N}$) of random couplings in region $\mu$
$\bm{m}^{\mu\nu}$ Vector with components $m_{i}^{\mu\nu}$; defines structured input pattern from region $\nu$ to neurons in region $\mu$ (Eq. 2)
$\bm{n}^{\mu\nu}$ Vector with components $n_{i}^{\mu\nu}$; defines structured readout pattern from neurons in region $\nu$ to region $\mu$ (Eq. 2)
DMFT variables
$\Delta^{\mu}(t,t^{\prime})$ Correlation function of preactivations in region $\mu$ (Eq. 5a)
$C^{\mu}(t,t^{\prime})$ Correlation function of activations in region $\mu$ (Eq. 5b)
$S^{\mu\nu}(t)$ Current from region $\nu$ to region $\mu$ at time $t$ (Eq. 6)
$H^{\mu\nu}(t)$ Drive to $S^{\mu\nu}(t)$ in the mean-field dynamics of the currents (Eq. 7)
$\psi^{\mu}(t)$ Neuronal gain in region $\mu$ at time $t$ (Eq. 8)
$A^{\mu}$ Sum of squared currents from all regions into region $\mu$
$S_{0}^{\mu\nu}$ Fixed-point value of inter-region current from region $\nu$ to region $\mu$
$\sigma^{\mu\nu}(t)$ Perturbation to the inter-region current from region $\nu$ to region $\mu$
$\hat{\Delta}^{\mu}(\tau)$ Normalized stationary correlation function (Eq. 21)
DMFT parameters
$T^{\mu\nu\rho}$ Normalized overlap between readout and input patterns, representing effective interaction from region $\rho$ to region $\mu$ through region $\nu$ (Eq. 4a)
$\hat{T}^{\mu\nu,\rho\sigma}$ Matrix form of $T^{\mu\nu\rho}$ (Eq. 11)
$U^{\mu\nu\rho}$ Overlap between input vectors in region $\mu$ originating from regions $\nu$ and $\rho$ (Eq. 4b)
$c^{\mu\nu}$ Symmetric parameterization of $T^{\mu\nu\rho}$ (Eq. 12)
$u^{\mu}$ Rank-one contribution to “rank-one plus diagonal” parameterization of $c^{\mu\nu}$ (Eq. 13)
$h^{\mu}$ Diagonal contribution to “rank-one plus diagonal” parameterization of $c^{\mu\nu}$ (Eq. 13)
$a^{\mu}$ Strength of direct self-interaction (Eq. 14)
$b^{\mu}$ Strength of indirect self-interaction (Eq. 14)
Table 1: Summary of notation.
4 Dynamical Mean-Field Theory (DMFT)
Mean-field theory is an analytical approach that describes large systems using a small set of summary statistics called order parameters. This method provides an exact description as $N\rightarrow\infty$ and a good approximation for large, finite $N$. Dynamical mean-field theory (DMFT) extends this concept by introducing time-dependent order parameters to capture the temporal evolution of activity [29, 47]. We now present the order parameters in the DMFT description of our multiregion network model and the equations governing their dynamics.
4.1 Order Parameters
Our multiregion model exhibits two types of dynamics: high-dimensional chaotic fluctuations from i.i.d. connectivity, and low-dimensional excitation within or between regions due to low-rank connectivity. These dynamics are described by distinct sets of order parameters.
High-dimensional fluctuations are characterized by correlation functions, which capture the temporal structure of chaotic fluctuations. For each region $\mu$, we define correlation functions for the (pre-)activations:
$\Delta^{\mu}(t,t^{\prime})$ $=\frac{1}{N}\sum_{i=1}^{N}x^{\mu}_{i}(t)x^{\mu}_{i}(t^{\prime}),$ (5a)
$C^{\mu}(t,t^{\prime})$ $=\frac{1}{N}\sum_{i=1}^{N}\phi^{\mu}_{i}(t)\phi^{\mu}_{i}(t^{\prime}).$ (5b)
Low-dimensional signal transmission within and between regions is described by currents, following the terminology of Perich et al. [12]. These currents are consolidated in the matrix $S^{\mu\nu}(t)$, defined by:
$\frac{dS^{\mu\nu}(t)}{dt}=-S^{\mu\nu}(t)+\frac{1}{N}\sum_{i=1}^{N}n^{\mu\nu}_{i}\phi_{i}^{\nu}(t).$ (6)
The current $S^{\mu\nu}(t)$ represents the activity in region $\nu$ that is transmitted to region $\mu$ (plus a low-pass filter).
4.2 Routing and Non-Routing Regions
The current matrix provides crucial information about activity flow between regions. We classify regions as routing or non-routing based on their role in signal transmission. We say that a region $\nu$ is routing if it transmits signals between other regions, indicated by at least one nonzero off-diagonal element in the $\nu$-th column of the current matrix, $S^{:,\nu}(t)$; and at least one nonzero off-diagonal element in the $\nu$-th row, $S^{\nu,:}(t)$. In contrast, we say that a region $\nu$ is non-routing if all elements of its corresponding row and column in the current matrix are zero, except possibly for the diagonal element, $S^{\nu\nu}(t)$.
As we will demonstrate through exact solutions of the DMFT equations, a region may become non-routing when its own activity is too strong, preventing signal flow. One way for this to occur is if the region’s activity aligns with its internal structured connectivity, resulting in a nonzero diagonal element, $S^{\nu\nu}\neq 0$.
Experimentally, routing of this type could be detected through analyses similar to those used by Semedo et al. [39]. By computing the communication subspace for a source region during spontaneous activity, one could see how activity patterns line up with that subspace during a task; the overlapping activity would be the routed signal.
4.3 Dynamical Mean-Field Equations
In the mean-field picture, currents interact according to:
$\frac{dS^{\mu\nu}(t)}{dt}$ $=-S^{\mu\nu}(t)+H^{\mu\nu}(t),\text{ where}$ (7a)
$H^{\mu\nu}(t)$ $=\psi^{\nu}(t)\sum_{\rho=1}^{R}{T}^{\mu\nu\rho}S^{\nu\rho}(t),$ (7b)
where $\psi^{\nu}(t)=\psi(\Delta^{\nu}(t,t))$ is the average gain of neurons in region $\nu$. The function $\psi(\Delta)$ performs a Gaussian average:
$\psi(\Delta)=\left\langle\phi^{\prime}(x)\right\rangle_{x},$ (8)
where $x\sim\mathcal{N}(0,\Delta)$. Thus, while standard neural networks have a vector dynamics shaped by a matrix, in our framework, region-to-region interactions, defined by the current order parameters, have a matrix dynamics shaped by a third-order tensor. Meanwhile, $\Delta^{\mu}(t,t^{\prime})$ satisfies:
$\left(1+\frac{d}{dt}\right)\left(1+\frac{d}{dt^{\prime}}\right)\Delta^{\mu}(t,t^{\prime})=(g^{\mu})^{2}C^{\mu}(t,t^{\prime})+\sum_{\nu,\rho=1}^{R}U^{\mu\nu\rho}H^{\mu\nu}(t)H^{\mu\rho}(t^{\prime}),$ (9)
@@ 4.3 Dynamical Mean-Field Equations
These equations are closed by expressing $C^{\mu}(t,t^{\prime})$ in terms of $\Delta^{\mu}(t,t^{\prime})$ via $C^{\mu}(t,t^{\prime})=C(\Delta(t,t^{\prime}),\Delta(t,t),\Delta(t^{\prime},t^{\prime}))$, where $C(\Delta_{12},\Delta_{11},\Delta_{22})$ propagates preactivation correlations to activation correlations:
$C(\Delta_{12},\Delta_{11},\Delta_{22})=\left\langle\phi(x_{1})\phi(x_{2})\right\rangle_{x_{1},x_{2}},$ (10)
where $(x_{1},x_{2})\sim\mathcal{N}\left(\bm{0},\bm{\Delta}\right)$. $\psi(\Delta)$ and $C(\Delta_{12},\Delta_{11},\Delta_{22})$ can be evaluated analytically (SI Appendix).
Thus, the DMFT provides a set of deterministic, causal dynamic equations for the region-specific two-point functions and currents. While their derivation is relatively straightforward, solving them analytically is challenging due to their nonlinear and time-dependent structure, as well as the tensorial form of the interactions. In the next section, we show that by assuming certain symmetry properties of $T^{\mu\nu\rho}$, we can, remarkably, derive a rich and instructive class of time-independent and time-dependent solutions.
@@
For the remainder of the paper, we assume $U^{\mu\nu\rho}=\delta^{\nu\rho}$ for all $\mu$, focusing on the role of $T^{\mu\nu\rho}$. Geometrically, this means that inputs from other regions into a target region $\mu$ are organized in orthogonal subspaces.
5 Symmetric Effective Interactions and Fixed Points
@@
We now set out to derive exact solutions to the DMFT equations. In general, to simplify the analysis of many-body interactions, a natural choice is to assume symmetry. In standard neural networks, symmetric interactions ensure that the system converges to fixed points, precluding limit cycles and chaos. However, enforcing symmetry in the DMFT system is challenging because the effective interactions among the currents form a third-order tensor, $T^{\mu\nu\rho}$.
To clarify the structure of the interactions between currents in the DMFT, we rewrite the right-hand side of the current dynamics as $\psi^{\nu}(t)\sum_{\rho,\sigma=1}^{R}\hat{T}^{\mu\nu,\rho\sigma}S^{\rho\sigma}(t),$ where
$\hat{T}^{\mu\nu,\rho\sigma}=\delta^{\nu\rho}T^{\mu\nu\sigma}$ (11)
is a $R^{2}$-by-$R^{2}$ dynamics matrix governing the linearized interaction of the $R^{2}$ currents (its spectrum is closely related to that of $J^{\mu\nu}_{ij}$; SI Appendix). We expect $\hat{T}^{\mu\nu,\rho\sigma}$ to influence the current dynamics similarly to how the synaptic weight matrix shapes neuronal dynamics in a standard neural network. Thus, a natural choice is to impose symmetry on the matrix $\hat{T}^{\mu\nu,\rho\sigma}$, i.e., $\hat{T}^{\mu\nu,\rho\sigma}=\hat{T}^{\rho\sigma,\mu\nu}$. This reduces the number of free parameters from $\mathcal{O}(R^{3})$ to $\mathcal{O}(R^{2})$ by requiring
$T^{\mu\nu\rho}=\delta^{\mu\rho}c^{\mu\nu},\text{ where }c^{\mu\nu}=c^{\nu\mu}.$ (12)
The presence of $\delta^{\mu\rho}$ in $T^{\mu\nu\rho}$ implies that each region $\mu$ interacts either directly with itself ($\mu=\nu$) or indirectly with itself through an intermediate region, $\nu$ ($\mu\neq\nu$). Moreover, the symmetry of $c_{\mu\nu}$ implies that the coupling through which region $\mu$ interacts with itself via region $\nu$ is equivalent to that for region $\nu$ interacting with itself via region $\mu$. This is illustrated in Fig. 2a.
To make analytical progress, we further constrain the symmetric matrix $c^{\mu\nu}$ to have a “rank-one plus diagonal” form, with only $\mathcal{O}(R)$ parameters,
@@
where $u^{\mu}$ and $h^{\mu}$ are arbitrary vectors. This form provides a minimal setting in which one has independent control over the strength of direct versus indirect self-interactions, which are captured by the quantities
$a^{\mu}=(u^{\mu})^{2}+h^{\mu},\text{ and }b^{\mu}=(u^{\mu})^{2},$ (14)
respectively. If $b^{\mu}=0$, region $\mu$ is not connected to the rest of the network, and its dynamical repertoire is that of a rank-one network with disorder, studied in [26].
5.1 Disorder-Free Case
We begin by examining the case without disorder in connectivity: $g^{\mu}=0$ for all $\mu$. Symmetric interactions typically lead to fixed points, which we find to be the case here (although we were unable to derive a global Lyapunov function). For the parameterization of $T^{\mu\nu\rho}$ defined above, the fixed points $S^{\mu\nu}_{0}$ of the currents satisfy: